# MIRA: Medical Time Series Foundation Model for Real-World Health Data

### v5

@@ Appendix A Further Related Work
(i) Encoder-only models. Moirai is trained on the LOTSA dataset comprising 27B time points, with model sizes up to 311M parameters [28]. MOMENT, based on the T5 architecture, is pretrained on the Time-series Pile dataset containing approximately 1B time points, reaching up to 385M parameters [84].
(ii) Encoder-decoder models, exemplified by Chronos [33], which tokenizes time series data via scaling and quantization, training on both public and synthetic datasets,offers pre-trained models at multiple scales, with the largest containing up to 710M parameters.
(iii) Decoder-only models. TimesFM is trained on a corpus of 100B time points, with model sizes up to 500M parameters [30]. Lag-Llama focuses on univariate probabilistic forecasting, utilizing a decoder-only Transformer architecture with up to 200M parameters [87]. Timer is a generative pre-trained Transformer model designed for large-scale time series modeling, with a base version containing 84M parameters and pre-trained on 260B time points [86].
(iv) Mixture-of-Experts architectures. Recent models adopt sparse Mixture-of-Experts (MoE) architectures to enhance scalability and efficiency. Time-MoE [29] scales to 2.4B parameters with only a few experts activated per input, while Moirai-MoE [31] achieves token-level specialization without frequency heuristics, improving adaptability and inference cost.
@@
(Ours) (2025) (2024) (2024a) (2024) (2024) (2024) (2024) (2024d) (2023) (2023)
@@
Appendix B Time Quantization Procedure
To ensure numerical stability and strictly increasing temporal inputs, we adopt a robust quantization strategy for irregular timestamps. The goal is to discretize time points with a suitable resolution while guaranteeing uniqueness. This is particularly important when encoding timestamps for attention-based models, where duplicate or non-monotonic inputs can lead to numerical or semantic inconsistencies.
The procedure begins with a predefined initial resolution (e.g., 1.0) and iteratively reduces it by a shrink factor (e.g., 10) until all quantized time points become unique. In each iteration, the algorithm rounds the original timestamps to the current resolution and checks for duplicates. If uniqueness is achieved, the quantized sequence is returned. If, after a fixed number of iterations, uniqueness cannot be guaranteed, a fallback mechanism is triggered. In this case, small additive jitter is applied to break ties while preserving temporal ordering. The jitter is dynamically determined based on the magnitude of the input values, ensuring that the perturbation is both numerically safe and practically negligible. Finally, the adjusted timestamps are cast to 32-bit floating-point precision and verified again for uniqueness. If rare collisions still exist due to floating-point rounding, a second corrective jitter step is applied [90, 91].
@@ Appendix D Neural ODE
For end-to-end training, gradients of a loss function $\mathcal{L}$ with respect to the ODE parameters $\theta_{\text{ODE}}$ and the initial state $h(t_{N})$ are efficiently computed using the adjoint sensitivity method [40]. This involves solving a backward-in-time ODE for the adjoint state $a(s)=\partial\mathcal{L}/\partial h(s)$:
@@ Appendix E Pretraining Datasets
(2023) (2016) (2020) (2000) (2020)
@@
MIMIC-III [73] is a widely used publicly available database containing de-identified health records of over 40,000 patients admitted to intensive care units (ICUs) at the Beth Israel Deaconess Medical Center between 2001 and 2012. It includes multivariate time series of vital signs, laboratory test results, medication administrations, and clinical notes. The data exhibit irregular sampling patterns and substantial missingness, making them suitable for evaluating temporal robustness.
MIMIC-IV [74] is the successor to MIMIC-III, covering ICU and emergency department admissions from 2008 to 2019. It offers expanded variable coverage, improved data standardization, and higher temporal resolution. Time series are recorded with greater granularity across a broader range of hospital units, including both adult and pediatric care. Like MIMIC-III, it reflects realistic clinical irregularity but with enhanced data fidelity and scale.
PTB-XL [75] is a 12-lead ECG dataset sampled at 100 Hz, containing over 20,000 clinical recordings annotated with diagnostic statements. It enables robust modeling of waveform-based cardiovascular signals.
Sleep-EDF [76] contains overnight sleep EEG and respiration recordings annotated with sleep stages, collected under controlled conditions. The data are sampled at 100 Hz and exhibit moderate regularity with biologically driven transitions.
WAVES [77] is a pediatric waveform dataset that combines high-frequency physiological signals (ECG, PPG, respiration) from intensive care and operating room settings. It provides long continuous sequences with natural variability and noise.
@@ Appendix F Model Configurations and Training Detail
Informed by recent scaling studies [29], we design MIRA in three model scales: $MIRA_{small}$ (73M parameters), $MIRA_{base}$ (114M), and $MIRA_{large}$ (455M), enabling flexible deployment across different hardware constraints. All models are trained on up to eight NVIDIA 80GB A100 GPUs with a micro-batch size of 128 and a maximum sequence length of 512. We pre-trained one epoch with each training step processes approximately 65,000 time points. We consider forecast horizons of 24,32,48,64{24,32,48,64} for short-term and long-term evaluation. Following standard practice, we apply an auxiliary load balancing loss with weight $\alpha=0.02$ to encourage expert utilization.
@@ Appendix G Downstream Tasks and Datasets
CinC 2012[78] originates from the PhysioNet 2012 Challenge and contains multivariate time series from ICU patients. Each record includes asynchronously sampled clinical variables such as blood pressure, heart rate, and oxygen saturation, with irregular measurement frequencies and observation patterns. The forecasting task focuses on predicting future physiological values and deterioration risk based on past observations.
CinC 2019[79] Derived from the PhysioNet 2019 Challenge, this dataset includes time series recorded during in-hospital stays, including variables like vital signs, lab test results, and intervention indicators. The data are highly sparse and irregular, reflecting realistic clinical workflows. We focus on continuous value prediction using sliding window evaluation, emphasizing temporal robustness.
@@
MIT-BIH [80] This dataset provides annotated ECG waveform segments from patients with arrhythmias. The input is a regularly sampled 2-lead ECG sequence, and the forecasting task involves predicting future signal windows. We introduce synthetic missingness during evaluation to test robustness.
JHU COVID-19 Dataset [81] This dataset aggregates daily COVID-19 cases and deaths by region. While data are reported at regular intervals, inconsistencies and reporting delays motivate the use of masked evaluation. Forecasting focuses on short-term case count prediction. We using this dataset in a daily frequency
CDC-Illness [83] The CDC outpatient illness dataset tracks ILI (influenza-like illness) and other symptoms across U.S. reporting centers. Although data are uniformly weekly, we mask 30% of time points to simulate partial surveillance dropout.
Heart Rate This dataset, sourced from the Monash Time Series Extrinsic Regression Archive [82], contains photoplethysmography (PPG) sequences paired with continuous heart rate values as targets. Each instance represents a short PPG segment with the goal of predicting the corresponding heart rate, making it a benchmark for physiological signal regression tasks.
@@ Appendix H Baselines
Time-MoE [29]: A sparsely activated decoder-only Transformer model that incorporates a Mixture-of-Experts (MoE) architecture with token-level routing. Trained on over 300 billion real-world time points spanning nine domains, Time-MoE demonstrates strong zero-shot forecasting performance, especially in long-range and multi-resolution prediction. It uses autoregressive decoding and sliding window inference with shared expert regularization.
@@
Moirai [28]: A universal forecasting backbone that uses patch-wise tokenization and any-variate self-attention. It is trained on the LOTSA dataset comprising 27 billion time points and supports forecasting across arbitrary time steps, variable sets, and resolutions. Moirai also employs resolution-adaptive projection layers to accommodate different patch sizes during inference.
@@
Moirai-MoE [31]: An extension of Moirai that integrates a token-level MoE module into the decoder block. This allows for frequency-specific specialization without relying on handcrafted signal partitions. Moirai-MoE achieves improved generalization across domains with limited added compute, and supports both token-aware routing and auxiliary balancing loss.
@@
TimesFM [30]: A decoder-only Transformer developed by Google Research, pre-trained on 100 billion real-world time points from diverse sources including IoT, finance, and weather. It features autoregressive generation with fixed-length context windows and shows strong performance in forecasting across hundreds of benchmarks.
@@
Chronos [33]: A family of probabilistic time series foundation models that transform numerical sequences into quantized token representations. Chronos leverages discrete latent spaces and causal language modeling for forecasting and sampling. It supports both point and probabilistic prediction.
@@
Timer [86]: A generative time series language model trained on a diverse mix of real-world and synthetic data. It uses next-token prediction and masked time-step modeling for flexible downstream adaptation. While architecture and scale details are less publicly documented, it has demonstrated competitive zero-shot accuracy on common forecasting benchmarks.
@@
ContiFormer [19]: A continuous-time Transformer model that combines Neural ODEs with attention. It uses a NeuralODE kernel for modeling value transitions between observations and a standard self-attention block for relational reasoning. It is trained end-to-end on irregularly sampled series.
@@
T-PatchGNN [88]: A graph-based model that converts univariate time series into a graph of overlapping time patches, using GNNs to capture local and global dependencies. It is especially effective for sparse and low-signal series and supports continuous-time node embedding.
@@
Neural-CDE [89]: A continuous-time model based on controlled differential equations. It encodes the trajectory of observed data using a Neural CDE solver and predicts future values through learned hidden dynamics. Particularly suited for datasets with asynchronous observations.
@@
ODE-RNN [66]: Combines standard RNN encoders with ODE solvers to update latent states over irregular intervals. It supports interpolation between time steps and improves temporal continuity compared to discrete RNNs.
@@
Moment [84] presents open-source Transformer-based models pre-trained on the Time-series Pile, a large and diverse collection of public time series data. Employing a masked time series prediction task, MOMENT demonstrates strong performance across forecasting, classification, and anomaly detection tasks, particularly in low-resource settings .
@@
Lag-Llama [87] is a decoder-only Transformer model tailored for univariate probabilistic forecasting. By incorporating lagged inputs as covariates and pretraining on a diverse corpus of time series data, Lag-Llama exhibits robust zero-shot generalization and state-of-the-art performance upon fine-tuning on unseen datasets .
@@
TimeGPT [85] is introduced as the first foundation model for time series analysis, capable of generating accurate predictions across diverse datasets without additional training. Leveraging advancements in deep learning, TimeGPT’s zero-shot inference outperforms traditional statistical, machine learning, and deep learning methods in both performance and efficiency .

### v6

@@ Appendix A Further Related Work
(i) Encoder-only models. Moirai is trained on the LOTSA dataset comprising 27B time points, with model sizes up to 311M parameters [DBLP:conf/icml/WooLKXSS24]. MOMENT, based on the T5 architecture, is pretrained on the Time-series Pile dataset containing approximately 1B time points, reaching up to 385M parameters [DBLP:conf/icml/GoswamiSCCLD24].
(ii) Encoder-decoder models, exemplified by Chronos [DBLP:journals/corr/abs-2403-07815], which tokenizes time series data via scaling and quantization, training on both public and synthetic datasets,offers pre-trained models at multiple scales, with the largest containing up to 710M parameters.
(iii) Decoder-only models. TimesFM is trained on a corpus of 100B time points, with model sizes up to 500M parameters [DBLP:conf/icml/DasKSZ24]. Lag-Llama focuses on univariate probabilistic forecasting, utilizing a decoder-only Transformer architecture with up to 200M parameters [rasul2023lag]. Timer is a generative pre-trained Transformer model designed for large-scale time series modeling, with a base version containing 84M parameters and pre-trained on 260B time points [DBLP:conf/icml/LiuZLH0L24].
(iv) Mixture-of-Experts architectures. Recent models adopt sparse Mixture-of-Experts (MoE) architectures to enhance scalability and efficiency. Time-MoE [DBLP:journals/corr/abs-2409-16040] scales to 2.4B parameters with only a few experts activated per input, while Moirai-MoE [DBLP:journals/corr/abs-2410-10469] achieves token-level specialization without frequency heuristics, improving adaptability and inference cost.
@@
(Ours) (DBLP:journals/corr/abs-2502-00816) (DBLP:journals/corr/abs-2409-16040) (DBLP:journals/corr/abs-2410-10469) (DBLP:conf/icml/WooLKXSS24) (DBLP:conf/icml/DasKSZ24) (DBLP:conf/icml/GoswamiSCCLD24) (DBLP:journals/corr/abs-2403-07815) (DBLP:conf/icml/LiuZLH0L24) (rasul2023lag) (DBLP:journals/corr/abs-2310-03589)
@@
Appendix B Time Normalization
MIRA operates in continuous time and relies on a rotary positional encoding adapted to irregular timestamps (CT-RoPE). To ensure stable phase computation and consistent temporal scaling across training and autoregressive inference, we apply a deterministic min–max normalization to each full timestamp sequence.
Given a strictly increasing timestamp sequence
$\mathbf{t}=[t_{1},t_{2},\ldots,t_{T}],$ (13)
we compute
$t_{\min}=\min_{i}\,t_{i},\qquad t_{\max}=\max_{i}\,t_{i},$ (14)
and rescale every timestamp into a canonical range that aligns with the sequence length:
$\widehat{t}_{i}=\frac{t_{i}-t_{\min}}{\,t_{\max}-t_{\min}+\varepsilon\,}\,(T-1).$ (15)
where, (T1)(T-1) corresponds to the maximum index of a sequence of length $T$, ensuring that the normalized time domain matches the index range used in standard rotary encodings, while preserving the original temporal spacing in a continuous form. The small constant $\varepsilon$ prevents numerical instability when the timestamp range is narrow [wu2025generative, wu2025prompt].
@@ Appendix D Neural ODE
For end-to-end training, gradients of a loss function $\mathcal{L}$ with respect to the ODE parameters $\theta_{\text{ODE}}$ and the initial state $h(t_{N})$ are efficiently computed using the adjoint sensitivity method [chen2018neural]. This involves solving a backward-in-time ODE for the adjoint state $a(s)=\partial\mathcal{L}/\partial h(s)$:
@@ Appendix E Pretraining Datasets
(miller2023waves) (johnson2016mimic) (johnson2020mimic) (kemp2000analysis) (wagner2020ptb)
@@
MIMIC-III [johnson2016mimic] is a widely used publicly available database containing de-identified health records of over 40,000 patients admitted to intensive care units (ICUs) at the Beth Israel Deaconess Medical Center between 2001 and 2012. It includes multivariate time series of vital signs, laboratory test results, medication administrations, and clinical notes. The data exhibit irregular sampling patterns and substantial missingness, making them suitable for evaluating temporal robustness.
MIMIC-IV [johnson2020mimic] is the successor to MIMIC-III, covering ICU and emergency department admissions from 2008 to 2019. It offers expanded variable coverage, improved data standardization, and higher temporal resolution. Time series are recorded with greater granularity across a broader range of hospital units, including both adult and pediatric care. Like MIMIC-III, it reflects realistic clinical irregularity but with enhanced data fidelity and scale.
PTB-XL [wagner2020ptb] is a 12-lead ECG dataset sampled at 100 Hz, containing over 20,000 clinical recordings annotated with diagnostic statements. It enables robust modeling of waveform-based cardiovascular signals.
Sleep-EDF [kemp2000analysis] contains overnight sleep EEG and respiration recordings annotated with sleep stages, collected under controlled conditions. The data are sampled at 100 Hz and exhibit moderate regularity with biologically driven transitions.
WAVES [miller2023waves] is a pediatric waveform dataset that combines high-frequency physiological signals (ECG, PPG, respiration) from intensive care and operating room settings. It provides long continuous sequences with natural variability and noise.
@@ Appendix F Model Configurations and Training Detail
Informed by recent scaling studies [DBLP:journals/corr/abs-2409-16040], we design MIRA in three model scales: $MIRA_{small}$ (73M parameters), $MIRA_{base}$ (114M), and $MIRA_{large}$ (455M), enabling flexible deployment across different hardware constraints. All models are trained on up to eight NVIDIA 80GB A100 GPUs with a micro-batch size of 128 and a maximum sequence length of 512. We pre-trained one epoch with each training step processes approximately 65,000 time points. We consider forecast horizons of 24,32,48,64{24,32,48,64} for short-term and long-term evaluation. Following standard practice, we apply an auxiliary load balancing loss with weight $\alpha=0.02$ to encourage expert utilization.
@@ Appendix G Downstream Tasks and Datasets
CinC 2012[silva2012predicting] originates from the PhysioNet 2012 Challenge and contains multivariate time series from ICU patients. Each record includes asynchronously sampled clinical variables such as blood pressure, heart rate, and oxygen saturation, with irregular measurement frequencies and observation patterns. The forecasting task focuses on predicting future physiological values and deterioration risk based on past observations.
CinC 2019[reyna2020early] Derived from the PhysioNet 2019 Challenge, this dataset includes time series recorded during in-hospital stays, including variables like vital signs, lab test results, and intervention indicators. The data are highly sparse and irregular, reflecting realistic clinical workflows. We focus on continuous value prediction using sliding window evaluation, emphasizing temporal robustness.
@@
MIT-BIH [moody2001impact] This dataset provides annotated ECG waveform segments from patients with arrhythmias. The input is a regularly sampled 2-lead ECG sequence, and the forecasting task involves predicting future signal windows. We introduce synthetic missingness during evaluation to test robustness.
JHU COVID-19 Dataset [dong2022johns] This dataset aggregates daily COVID-19 cases and deaths by region. While data are reported at regular intervals, inconsistencies and reporting delays motivate the use of masked evaluation. Forecasting focuses on short-term case count prediction. We using this dataset in a daily frequency
CDC-Illness [DBLP:conf/iclr/WuHLZ0L23] The CDC outpatient illness dataset tracks ILI (influenza-like illness) and other symptoms across U.S. reporting centers. Although data are uniformly weekly, we mask 30% of time points to simulate partial surveillance dropout.
Heart Rate This dataset, sourced from the Monash Time Series Extrinsic Regression Archive [tan2020monash], contains photoplethysmography (PPG) sequences paired with continuous heart rate values as targets. Each instance represents a short PPG segment with the goal of predicting the corresponding heart rate, making it a benchmark for physiological signal regression tasks.
@@ Appendix H Baselines
Time-MoE [DBLP:journals/corr/abs-2409-16040]: A sparsely activated decoder-only Transformer model that incorporates a Mixture-of-Experts (MoE) architecture with token-level routing. Trained on over 300 billion real-world time points spanning nine domains, Time-MoE demonstrates strong zero-shot forecasting performance, especially in long-range and multi-resolution prediction. It uses autoregressive decoding and sliding window inference with shared expert regularization.
@@
Moirai [DBLP:conf/icml/WooLKXSS24]: A universal forecasting backbone that uses patch-wise tokenization and any-variate self-attention. It is trained on the LOTSA dataset comprising 27 billion time points and supports forecasting across arbitrary time steps, variable sets, and resolutions. Moirai also employs resolution-adaptive projection layers to accommodate different patch sizes during inference.
@@
Moirai-MoE [DBLP:journals/corr/abs-2410-10469]: An extension of Moirai that integrates a token-level MoE module into the decoder block. This allows for frequency-specific specialization without relying on handcrafted signal partitions. Moirai-MoE achieves improved generalization across domains with limited added compute, and supports both token-aware routing and auxiliary balancing loss.
@@
TimesFM [DBLP:conf/icml/DasKSZ24]: A decoder-only Transformer developed by Google Research, pre-trained on 100 billion real-world time points from diverse sources including IoT, finance, and weather. It features autoregressive generation with fixed-length context windows and shows strong performance in forecasting across hundreds of benchmarks.
@@
Chronos [DBLP:journals/corr/abs-2403-07815]: A family of probabilistic time series foundation models that transform numerical sequences into quantized token representations. Chronos leverages discrete latent spaces and causal language modeling for forecasting and sampling. It supports both point and probabilistic prediction.
@@
Timer [DBLP:conf/icml/LiuZLH0L24]: A generative time series language model trained on a diverse mix of real-world and synthetic data. It uses next-token prediction and masked time-step modeling for flexible downstream adaptation. While architecture and scale details are less publicly documented, it has demonstrated competitive zero-shot accuracy on common forecasting benchmarks.
@@
ContiFormer [DBLP:conf/nips/ChenRWFSL23]: A continuous-time Transformer model that combines Neural ODEs with attention. It uses a NeuralODE kernel for modeling value transitions between observations and a standard self-attention block for relational reasoning. It is trained end-to-end on irregularly sampled series.
@@
T-PatchGNN [DBLP:conf/icml/ZhangYL0024]: A graph-based model that converts univariate time series into a graph of overlapping time patches, using GNNs to capture local and global dependencies. It is especially effective for sparse and low-signal series and supports continuous-time node embedding.
@@
Neural-CDE [DBLP:conf/nips/RubanovaCD19]: A continuous-time model based on controlled differential equations. It encodes the trajectory of observed data using a Neural CDE solver and predicts future values through learned hidden dynamics. Particularly suited for datasets with asynchronous observations.
@@
ODE-RNN [DBLP:journals/corr/abs-1907-03907]: Combines standard RNN encoders with ODE solvers to update latent states over irregular intervals. It supports interpolation between time steps and improves temporal continuity compared to discrete RNNs.
@@
Moment [DBLP:conf/icml/GoswamiSCCLD24] presents open-source Transformer-based models pre-trained on the Time-series Pile, a large and diverse collection of public time series data. Employing a masked time series prediction task, MOMENT demonstrates strong performance across forecasting, classification, and anomaly detection tasks, particularly in low-resource settings .
@@
Lag-Llama [rasul2023lag] is a decoder-only Transformer model tailored for univariate probabilistic forecasting. By incorporating lagged inputs as covariates and pretraining on a diverse corpus of time series data, Lag-Llama exhibits robust zero-shot generalization and state-of-the-art performance upon fine-tuning on unseen datasets .
@@
TimeGPT [DBLP:journals/corr/abs-2310-03589] is introduced as the first foundation model for time series analysis, capable of generating accurate predictions across diverse datasets without additional training. Leveraging advancements in deep learning, TimeGPT’s zero-shot inference outperforms traditional statistical, machine learning, and deep learning methods in both performance and efficiency .